# A Bayesian hierarchical model for methane emission source apportionment

### v1

@@ 4.3 Comparison to alternative methods
We analyze the METEC experiment using two basic statistical techniques (OLS and LASSO) and three state-of-the-art inversion methods from the literature, with results summarized in Table 3. We select two single-source methods to compare against (Cartwright et al., 2019; Daniels et al., 2024a), both of which assume that only one source is emitting at a time. To apply them to the METEC data, we use the localization procedure from Daniels et al. (2024a) to estimate a most likely source for each inversion window before the inversion is performed. We select one multisource method to compare against (Weidmann et al., 2022), which can either estimate fluxes on a fixed grid or identify emitting sources using a reversible jump MCMC sampler. To provide a fair comparison, we use this method with a fixed grid that is constrained to the locations of the five sources on the METEC site, as the MDLQ model specifies these sources a priori. Finally, we also compare to a variant of the MDLQ model that assumes independent Laplace errors. Through basic model diagnostics, we found that errors from the MDLQ model often have heavier tails than a Gaussian. These large errors are likely from instances in which a spike in the concentration predictions is not present in the observations, or visa versa. We believed that a Laplace error model might be better able to capture large errors and hence result in better coverage. Section S9 of the SI fully defines this model variant. As none of these alternative methods discuss how to best use their inversion method in practice (rather than on releases with known start and end times), we use the moving time window approach discussed in Section 3.3 for all methods.
@@ methods
The MDLQ model has the smallest error in the site-level inventory and in the individual site-level emission rate estimates. No one method has the smallest error across the five source-level inventories. Weidmann et al. (2022) provides the best Tank and East Wellhead inventory, Cartwright et al. (2019) provides the best West Wellhead and West Separator inventory, and OLS provides the best East Separator inventory. Weidmann et al. (2022) has the smallest IQR. The single-source inversion method from Daniels et al. (2024a) has the best coverage, followed by the MDLQ. The MDLQ model variant with Laplace errors has the best localization performance, closely followed by OLS. Note that for OLS, LASSO, and the method from Weidmann et al. (2022), we assume that a source is emitting if its rate estimate is above 0.1 kg/hr, which was hand-tuned to provide the best possible localization performance on these data.
@@ 5 Discussion
We have proposed a Bayesian hierarchical model for methane emission source apportionment on oil and gas sites, referred to as the MDLQ model. This model is able to estimate total emissions on a representative oil and gas site to within 0.1 metric tons over the course of a three month experiment, a percent error of just -1.0%.
@@
The primary statistical contribution of this work is the synthesis and extension of other source apportionment methods from the existing literature, such as the spike-and-slab prior from Weidmann et al. (2022), the autocorrelated error model from Ganesan et al. (2014), and the time-varying Gaussian puff atmospheric dispersion model from Jia et al. (2025). We extend these studies and synthesize their contributions in the MDLQ model. Our implementation of the model is able to achieve a 100% convergence rate due to several innovative steps of the Gibbs sampler that are described fully in Section S3 of the SI.
@@
The primary contribution of this work to the field of methane emissions mitigation is twofold. First, we provide the most accurate CMS-based inversion method for creating site-level emissions inventories to date. Site-level inventories are an important tool for emissions accounting and mitigation in US regulations (EPA, 2024), EU import standards (Council of European Union, 2023), and voluntary emission management protocols (OGMP 2.0, 2024). The inversion performance of many commercially available CMS solutions is lacking (Ilonze et al., 2024; Chen et al., 2024; Daniels et al., 2025), and we hope to elevate all solutions operating in this space by making the MDLQ model implementation fully open source. Second, the METEC experiment provides a more robust evaluation of source apportionment methods for oil and gas sites than other experiments in the literature, as it contains about 10 times more releases and uses more realistic emission characteristics. As such, we believe that the results in this paper are more likely to generalize to real oil and gas sites, and by making all data from this study open source, we hope to provide a robust test bed for future modeling efforts.
@@
We conclude by discussing two limitations of the MDLQ model and future lines of research that may address them. First, the MDLQ model does not learn any parameters of the forward atmospheric dispersion model as is done in Cartwright et al. (2019) or Newman et al. (2024) and instead prescribes them based on atmospheric stability class. This may contribute to the large variability in inferred emission rates, as the Pasquill-Gifford-Turner dispersion scheme can overestimate or underestimate true atmospheric dispersion. Developing a faster implementation of the Gaussian puff model (or another fast solution to the advection-diffusion equation) may make it feasible to sample dispersion parameters within the MCMC without having to use the lower fidelity Gaussian plume model (Jia et al., 2025). Second, errors in the source-level inventories range from -38.8% to 61.4% when using all available data, but are significantly reduced when using just the inversion windows that have at least one sensor downwind of all sources for the entire window. With this in mind, we will investigate an adaptive windowing approach where the exact boundaries of each inversion window are based on the information content of the CMS data rather than being fixed at 30-minute intervals. Additionally, using a probabilistic sampling scheme similar to the one proposed in Daniels et al. (2024b) may provide a better way to extrapolate emission rate estimates from windows with full information to the desired temporal range (e.g., an annual inventory for regulatory reporting). We believe that these efforts would improve source-level inventories and alerts derived from the MDLQ model.
Code and data availability
All code and data required to reproduce the results in this manuscript will be made open source upon publication. Anonymized code and data will be made available to referees during the review process.

### v2

@@ 4.3 Comparison to alternative methods
We analyze the METEC experiment using two other statistical techniques (OLS and LASSO) and three state-of-the-art inversion methods from the literature, with results summarized in Table 3. We select two single-source methods to compare against (Cartwright et al., 2019; Daniels et al., 2024a), both of which assume that only one source is emitting at a time. To apply them to the METEC data, we use the localization procedure from Daniels et al. (2024a) to estimate a most likely source for each inversion window before the inversion is performed. It is important to benchmark the performance of single-source methods on multisource releases, despite this being a clear violation of their core assumption, because these methods are still widely used in practice by sensor technology vendors. As such, this comparison provides a baseline level of apportionment accuracy that should be expected on real oil and gas sites from single-source inversion systems.
We also compare to pyELQ, a multisource inverse model based on the method from Weidmann et al. (2022) and IJzermans et al. (2024). The pyELQ model can estimate methane fluxes on a fixed grid or identify emitting sources using a reversible jump MCMC sampler. To provide a fair comparison, we run pyELQ with a fixed grid that is constrained to the locations of the five sources on the METEC site, as the MDLQ model specifies these sources a priori. Finally, we also compare to a variant of the MDLQ model that assumes independent Laplace errors. Through basic model diagnostics, we found that errors from the MDLQ model often have heavier tails than a Gaussian. These large errors are likely from instances in which a spike in the concentration predictions is not present in the observations, or vice versa. We believed that a Laplace error model might be better able to capture large errors and hence result in better coverage. Section S9 of the SI fully defines this model variant. As none of these alternative methods discuss how to best use their inversion method in practice (rather than on releases with known start and end times), we use the moving time window approach discussed in Section 3.3 for all methods.
The MDLQ model has the smallest error in the site-level inventory and in the individual site-level emission rate estimates. No one method has the smallest error across all five source-level inventories. The pyELQ model produces the best Tank inventory, LASSO produces the best West Wellhead, West Separator, and East Wellhead inventories, and OLS produces the best East Separator inventory. The MDLQ model variant with Laplace errors has the smallest IQR. The MDLQ model has the best site-level coverage of the 95% credible intervals. OLS has the best localization performance, followed by the Laplace model variant. Note that for OLS, LASSO, and pyELQ, we assume that a source is emitting if its rate estimate is above 0.1 kg/hr, which was hand-tuned to provide the best possible localization performance on these data.
@@ 5 Discussion
We have proposed a Bayesian hierarchical model for methane emission source apportionment on oil and gas sites, referred to as the MDLQ model. This model is able to estimate total emissions on a representative oil and gas site to within 0.11 metric tons over the course of a three month experiment, a percent error of just 4.5%.
@@
The primary statistical contribution of this work is the synthesis and extension of other source apportionment methods from the existing literature, such as the spike-and-slab prior from Weidmann et al. (2022), the autocorrelated error model from Ganesan et al. (2014), and the time-varying Gaussian puff atmospheric dispersion model from Jia et al. (2025). We extend these studies and synthesize their contributions in the MDLQ model.
@@
The contribution of this work to the field of methane emissions mitigation is twofold. First, we provide the most accurate method for constructing site-level emissions inventories with in situ sensors to date. Site-level inventories are an important tool for emissions accounting and mitigation in US regulations (EPA, 2024), EU import standards (Council of European Union, 2023), and voluntary emission management protocols (OGMP 2.0, 2024). The inversion performance of many commercially available CMS solutions is lacking (Ilonze et al., 2024; Chen et al., 2024; Daniels et al., 2025), and we hope to elevate all solutions for this problem by making the MDLQ model open source. Second, the METEC experiment provides a more robust evaluation of source apportionment methods for oil and gas sites than other experiments in the literature, as it contains about 10 times more releases and uses more realistic emission characteristics. As such, we believe that the results in this paper are more likely to generalize to real oil and gas sites.
@@
We conclude by discussing several limitations of the MDLQ model and METEC experiment and how they may be addressed in future work. First, the MDLQ model does not estimate or adjust any parameters of the forward atmospheric dispersion model as is done in Cartwright et al. (2019) or Newman et al. (2024) and instead prescribes them based on atmospheric stability class. This may contribute to the large variability in inferred emission rates, as the Pasquill-Gifford-Turner dispersion scheme can overestimate or underestimate true atmospheric dispersion. Developing a faster implementation of the Gaussian puff model (or another fast solution to the advection-diffusion equation) may make it feasible to sample dispersion parameters within the MCMC without having to use the lower fidelity Gaussian plume model.
Second, the METEC facility is designed to emulate relatively small oil and gas production sites with spatially distinct equipment groups. Other oil and gas facilities, such as midstream compressor stations, can be much larger with centrally housed equipment and overlapping emission sources. The MDLQ model may be able to apportion emissions between broad segments of these sites rather than specific equipment groups, but more testing is needed to assess this potential use case. More broadly, the MDLQ model is currently focused on apportionment problems where the number and location of potential sources is known. Oil and gas production sites are an important example of this type of apportionment problem, but they exist in other application domains as well, such as engineered landfills and carbon capture and storage facilities. Future work will extend the MDLQ model to apportionment problems where the number of potential sources is unknown. For example, the MDLQ model could be used to regularize a grid of emission sources using the spike-and-slab prior, similar to the method in Weidmann et al. (2022), or it could be used to refine a coarser estimate of source location. For the latter, the single-source localization procedure from Daniels et al. (2024a) could be iteratively applied to a grid of emission sources to identify the subset whose emissions (when forward simulated) exhibit strong correlation with observed concentrations; the MDLQ model could then further refine this subset.
Third, errors in the source-level inventories from the MDLQ model range from -37.7% to 81.8% when using all available data, but are significantly reduced when using just the inversion windows that have at least one sensor downwind of all sources for the entire window. With this in mind, we will investigate an adaptive windowing approach where the exact boundaries of each window are based on the information content of the CMS data rather than being fixed at 30-minute intervals. Additionally, using a probabilistic sampling scheme similar to the one proposed in Daniels et al. (2024b) may provide a better way to extrapolate emission rate estimates from windows with full information to the desired temporal range (e.g., an annual inventory for regulatory reporting). We believe that these efforts would improve source-level inventories and alerts derived from the MDLQ model.
6 Significance Statement
Efforts to reduce methane emissions from the oil and gas sector are increasingly based on measurements of ambient methane concentrations, but using these data to apportion emissions between sources on individual sites remains a challenge. This study pairs a new method for source apportionment with a controlled release dataset for model evaluation that accurately mimics real-world emission characteristics. By accounting for emission intermittency and autocorrelation through time, the model demonstrates state-of-the-art apportionment accuracy and uncertainty quantification. By making the model open-source, it can be adopted by oil and gas companies to prioritize their mitigation efforts or by regulators to verify their long-term emission reduction requirements. More broadly, the model is applicable to other source apportionment problems where the number and location of potential sources are known, extending its relevance beyond oil and gas methane emissions.
{acks}
[Acknowledgments] The authors would like to thank Cal Okenberg for identifying the TruncatedNormal package in R for use in the MCMC sampler and for reviewing the manuscript. The authors would also like to thank Troy Sorensen and Spencer Kidd for testing the MDLQ algorithm on multiple oil and gas sites and for reviewing the manuscript.